# RanPAC: Random Projections and Pre-trained Models for Continual Learning

### v1

@@ F.5 Scaling with projection size
Table A5: Scaling with $M$ for split CIFAR100. The table shows final average accuracy for $T=10$ and no Phase 1, with $\lambda=100$ as a constant for all $M$. The backbone network is ViT-B/16 trained on ImageNet21K.
@@ F.7 Experiments with ResNets
Unlike prompting strategies, our approach works with any feature extractor, e.g. both pre-trained transformer networks and pre-trained convolutional neural networks. To illustrate this, Table A6 and Table A7 shows results for ResNet50 and ResNet 152, respectively, pre-trained on ImageNet. We used $T=10$ tasks. Although this is different to the $T=20$ tasks used for ResNets by zhou2023revisiting, the accuracies we report for NCM are very comparable to those in zhou2023revisiting. As with results for pre-trained ViT-B/16 networks, the use of random projections and second-order statistics both provide significant performance boosts compared with NCM alone. We do not use Phase 1 of Algorithm 1 here but as shown by panos2023session; zhou2023revisiting, this is feasible for diverse PETL methods for convolutional neural networks. Interestingly, ResNet152 produces reduced error rates compared to ResNet50 on CUB, OmniBenchmark and Cars. It is possible that this is because the optimal value of $\lambda$ was outside the range considered in our experiments.
@@
Phase 2 $73.5\%$ $56.4\%$ $31.2\%$ $66.8\%$ $66.6\%$ $88.8\%$ $48.0$%
Phase 2 (No RPs) $69.0\%$ $55.3\%$ $31.1\%$ $58.7\%$ $58.7\%$ $87.7\%$ $42.5$%
NCM $56.8\%$ $43.0\%$ $22.3\%$ $50.6\%$ $55.0\%$ $81.2\%$ $26.5$%
Table A6: Results for ResNet50. Results are final Average Accuracies for $T=10$ tasks.
@@ F.8 Experiments with CLIP vision model
To further verify the general applicability of our method, we show results for a CLIP Radford vision model as the backbone pre-trained model in Table A8. The same general trend as for pre-trained ViT-B/16 models and ResNets can be seen, where use of RPs (Phase 2 in Algorithm 1) produces better accuracies than NCM alone. Interestingly, the results for Cars is substantially better with the CLIP vision backbone than for ViT-B/16 networks, whereas the opposite is true for the other datasets. It is possible that data from a very similar domain as the Cars dataset was included in training of the CLIP vision model.
@@
Phase 2 $76.6\%$ $59.3\%$ $35.0\%$ $65.3\%$ $62.7\%$ $89.1\%$ $40.0$%
Phase 2 (No RPs) $73.6\%$ $57.3\%$ $34.6\%$ $57.9\%$ $56.3\%$ $87.8\%$ $40.3$%
NCM $64.6\%$ $47.3\%$ $26.7\%$ $49.4\%$ $54.7\%$ $82.0\%$ $25.9$%
Table A7: Results for ResNet152. Results are final Average Accuracies for $T=10$ tasks.
Method CIFAR100 IN-R IN-A CUB OB VTAB Cars
Phase 2 $73.7\%$ $73.9\%$ $28.9\%$ $71.4\%$ $69.6\%$ $89.4\%$ $87.3$%
Phase 2 (No RPs) $68.5\%$ $65.3\%$ $25.7\%$ $65.2\%$ $60.0\%$ $87.9\%$ $82.3$%
NCM $59.1\%$ $65.9\%$ $21.9\%$ $63.2\%$ $59.6\%$ $73.9\%$ $82.1$%
Table A8: Results for CLIP-ViT-B/32. Results are final Average Accuracies for $T=10$ tasks. We used the pre-trained weights from OpenCLIP ilharco_gabriel_2021_5143773 (CLIP-ViT-B-32-laion2B-s34B-b79K).

### v2

@@ F.5 Scaling with projection size
Table A5: Scaling with $M$ for split CIFAR100. The table shows final average accuracy for $T=10$ and no Phase 1, with $\lambda=100$ as a constant for all $M$, using ViT-B/16 trained on ImageNet21K.
@@ F.7 Experiments with ResNets
Unlike prompting strategies, our approach works with any feature extractor, e.g. both pre-trained transformer networks and pre-trained convolutional neural networks. To illustrate this, Table A6 and Table A7 shows results for ResNet50 and ResNet 152, respectively, pre-trained on ImageNet. We used $T=10$ tasks (except for VTAB, which is $T=5$ tasks). Although this is different to the $T=20$ tasks used for ResNets by zhou2023revisiting, the accuracies we report for NCM are very comparable to those in zhou2023revisiting. As with results for pre-trained ViT-B/16 networks, the use of random projections and second-order statistics both provide significant performance boosts compared with NCM alone. We do not use Phase 1 of Algorithm 1 here but as shown by panos2023session, zhou2023revisiting, this is feasible for diverse PETL methods for convolutional neural networks. Interestingly, ResNet152 with RP produces reduced accuracies compared to ResNet50 on CUB, Omnibenchmark and VTAB. It is possible that this would be remedied by seeking an optimal value of $M$, whereas for simplicity we chose $M=10000$. Note that unlike the pre-trained ViT-B/16 model, the pre-trained ResNets require preprocessing to normalize the input images. We found, however, that this is best removed for VTAB.
@@
Phase 2 $77.5\%$ $56.8\%$ $36.1\%$ $70.9\%$ $68.0\%$ $88.8\%$ $48.6$%
Phase 2 (No RPs) $73.3\%$ $55.0\%$ $37.0\%$ $62.6\%$ $60.27\%$ $87.7\%$ $42.7$%
NCM $61.5\%$ $43.1\%$ $25.8\%$ $55.3\%$ $57.4\%$ $81.2\%$ $26.2$%
Table A6: Results for ResNet50. Results are final Average Accuracies for $T=10$ tasks, except VTAB which is $T=5$ tasks.
Method CIFAR100 IN-R IN-A CUB OB VTAB Cars
Phase 2 $79.7\%$ $59.0\%$ $41.4\%$ $66.6\%$ $65.8\%$ $88.0\%$ $45.1$%
Phase 2 (No RPs) $77.8\%$ $57.1\%$ $40.4\%$ $59.4\%$ $58.0\%$ $87.8\%$ $38.8$%
NCM $70.1\%$ $47.6\%$ $31.9\%$ $50.4\%$ $56.1\%$ $82.0\%$ $25.1$%
Table A7: Results for ResNet152. Results are final Average Accuracies for $T=10$ tasks, except VTAB which is $T=5$ tasks.
@@ F.8 Experiments with CLIP vision model
To further verify the general applicability of our method, we show results for a CLIP Radford vision model as the backbone pre-trained model in Table A8. The same general trend as for pre-trained ViT-B/16 models and ResNets can be seen, where use of RPs (Phase 2 in Algorithm 1) produces better accuracies than NCM alone. Interestingly, the results for Cars is substantially better with the CLIP vision backbone than for ViT-B/16 networks. It is possible that data from a very similar domain as the Cars dataset was included in training of the CLIP vision model. The CLIP result for Phase 2 only (see ablations in Table 1) is also better for Imagenet-R than for ViT/B-16, but for all other datasets, ViT/B-16 has higher accuracy. Note that unlike the pre-trained ViT-B/16 model, the pre-trained CLIP vision model requires preprocessing to normalize the input images.
@@
Phase 2 $85.0\%$ $76.5\%$ $40.2\%$ $79.4\%$ $75.1\%$ $90.5\%$ $90.5$%
Phase 2 (No RPs) $80.9\%$ $67.4\%$ $37.2\%$ $72.7\%$ $66.3\%$ $89.8\%$ $85.9$%
NCM $77.1\%$ $68.1\%$ $32.6\%$ $73.8\%$ $67.8\%$ $78.8\%$ $85.6$%
Table A8: Results for CLIP-ViT-B/16. Results are final Average Accuracies for $T=10$ tasks, except VTAB which is $T=5$ tasks. We used the pre-trained weights from OpenCLIP, finetuned on ImageNet-1K ilharco_gabriel_2021_5143773.